Circular No. 7460

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

S/1999 J 1 While examining measurements made by the University of Arizona’s

Spacewatch program in fields selected by J. V. Scotti for discovering outer satellites of Jupiter, T. B. Spahr, Center for Astrophysics, identified evident prediscovery observations of the object for which Spacewatch observations had been published on MPS 7418 under the designation 1999 UX_{18}. The slow motion and impossibility of fitting the observations with a heliocentric orbit suggested that the object might be a jovian satellite. R. S. McMillan, J. A. Larsen, J. Montani, A. E. Gleason and T. Gehrels were also involved with the observations and/or measurements. An unperturbed jovicentric solution by B. G. Marsden indeed proved the object to be a satellite, and this was used by G. V. Williams to obtain the perturbed solution below.

1999 UT  α_{2000}  δ_{2000}  V
Oct. 6.36219  2^{h}05^{m}27.^{s}71  +12^{o}28′47′′.8  19.9
6.38353  2 05 27.22  +12 28 44.8  20.2
6.40498  2 05 26.69  +12 28 41.9  20.5
12.36523  2 02 57.27  +12 14 55.4  20.2
12.40965  2 02 56.06  +12 14 49.6  20.7
19.33318  1 59 49.99  +11 57 15.0  20.5
19.35644  1 59 49.33  +11 57 11.3  20.7
19.37792  1 59 48.71  +11 57 07.9  20.6
30.33261  1 54 44.66  +11 27 38.6  21.2
30.35371  1 54 44.08  +11 27 34.9  21.0
30.37484  1 54 43.55  +11 27 31.5  20.7
Nov. 4.36652  1 52 28.07  +11 14 03.0  20.8
4.38952  1 52 27.39  +11 13 59.1  21.1
4.41569  1 52 26.69  +11 13 54.9  20.7

Epoch = 2000 Sept. 13.0 TT

T = 2001 Sept. 4.042 TT  ω = 53.^{o}891
e = 0.12971  Ω = 280.375 }2000.0
q = 0.14134 AU  i = 143.026

a = 0.16241 AU n^{o} = 0.46530 P = 774 days

2000 TT α_{2000} δ_{2000} Δ r_{J} ϵ Δα′′ Δδ′′
July 25 4^{h}11.^{m}92 +19^{o}26.^{′}8 5.640 0.184 57.^{o}8  16+ 2892−
Aug. 4 4 18.24 +19 38.1 5.503 0.184 65.8 529− 3234−
14 4 23.75 +19 46.3 5.355 0.184 74.1 1072− 3571−
24 4 28.33 +19 51.7 5.200 0.184 82.6 1608− 3902−
2000 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden